OpenStudy (anonymous):

(Definite intergral) (1/ x^2 (radx^2 +1))dx using trig functions

OpenStudy (amistre64):

what does the (x^2+1) part tell us about which sub we use?

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1}{x^2 \sqrt{x^2+1}} dx?\]

OpenStudy (anonymous):

that we can let u = tan u

OpenStudy (anonymous):

Yes

myininaya (myininaya):

close. let x=tan(u)

OpenStudy (anonymous):

ok, my apologies

myininaya (myininaya):

sorry doesn't cut it i was offended when u said let u=tan u lol

OpenStudy (anonymous):

lol

myininaya (myininaya):

ok so what is the next step?

OpenStudy (anonymous):

This is my first time doing a problem like this

myininaya (myininaya):

take derivative of both sides

OpenStudy (amistre64):

my comp froze :(

OpenStudy (anonymous):

i believe we differenate

myininaya (myininaya):

right! :)

OpenStudy (anonymous):

ok

myininaya (myininaya):

is this right if x=tan(u) then dx=sec^2(u) du ?

OpenStudy (anonymous):

I agree myininaya

myininaya (myininaya):

ok now lets make our substitutions in the integral

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1}{(\tan(u))^2 \sqrt{tan(u))^2+1}} \sec^2(u) du\]

myininaya (myininaya):

recall your trig identity \[\tan^2(u)+1=\sec^2(u)\]

OpenStudy (anonymous):

ok I do and I got that part :-)

myininaya (myininaya):

\[\int\limits_{}^{}\frac{\sec^2(u)}{\tan^2(u) \sqrt{\sec^2(u)}} du\]

myininaya (myininaya):

\[\int\limits_{}^{}\frac{\sec^2(u)}{\tan^2(u) \sec(u)}\]

myininaya (myininaya):

\[\int\limits_{}^{}\frac{\sec(u)}{\tan^2(u)} du\]

OpenStudy (anonymous):

alright I 'm following and not lost thus far

myininaya (myininaya):

\[\int\limits_{}^{}\sec(u) \cdot \frac{1}{\tan^2(u)} du=\int\limits_{}^{}\frac{1}{\cos(u)} \cdot \frac{\cos^2(u)}{\sin^2(u)} du\]

myininaya (myininaya):

\[\int\limits_{}^{}\frac{\cos(u)}{\sin^2(u)} du\]

myininaya (myininaya):

guess what? we are almost there

myininaya (myininaya):

do you see what to do next

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

let theta = sin u

myininaya (myininaya):

ok yes we can call the new substitution theta ! :)

OpenStudy (anonymous):

dtheta = cos u du

myininaya (myininaya):

incredible :)

OpenStudy (anonymous):

cool then it would be intergral 1/theta^2

myininaya (myininaya):

yes thats what we want to integrate

myininaya (myininaya):

oh you said integral gj

OpenStudy (anonymous):

:-)

myininaya (myininaya):

\[\frac{\theta^{-2+1}}{-2+1}+C=\frac{\theta^{-1}}{-1}+C=\frac{-1}{\theta}+C\]

myininaya (myininaya):

but what was theta?

myininaya (myininaya):

sin(u) right?

OpenStudy (anonymous):

sin u

OpenStudy (anonymous):

yes

myininaya (myininaya):

omg you are so great

myininaya (myininaya):

\[\frac{-1}{\sin(u)}+C\]

OpenStudy (anonymous):

lol why thanks lol

myininaya (myininaya):

last thing to do is rewrite in terms of x

OpenStudy (anonymous):

ok

myininaya (myininaya):

so remember we let x=tan(u) you might feel a need to draw a triangle thats what i would do